Grade 12General Physics

A uniform rope of length L lies in a lump on a frictionless table. a
very small piece at one end is held hanging down through a hole in the table just below the lump.
This system is released, and the rope slides down through the hole. What is the
speed of the other end of rope when it just leaves the table?

sarthak sharma
11 Years agoGrade 12

1 Answer

Mallikarjun Maram
11 Years ago
This can be solved by applying conservation of mechanical energy.  Let the table is at height H (>= L). Then
KEi = 0, PEi = mgH; KEf = (1/2)mv2, PEf = mg(H – L/2)
Substituting values in KEi + PEi = KEf + PEf and solving, we get v = sqrt(gL).
Answer will differ if H